OpenStudy (anonymous):

Let f(x) = x2 – 8x + 5. Find f(–1).

OpenStudy (anonymous):

\[f(x) = x^{2} - 8x + 5\]\[f(-1) = (-1)^{2} - 8(-1) + 5\]\[f(-1) = 1 + 8 + 5\]\[f(-1) = ?\]

OpenStudy (anonymous):

14

OpenStudy (anonymous):

Yup. \[\huge\color{red}{f(-1) = 14}\]